If a light with a frequency of radiation of 8×10^14 Hz is shone on metal and photoelectron are ejected with a maximum kinetic energy of 1.6×10^-19 J, what is work function of the material ?


1
Expert's answer
2021-06-01T18:06:41-0400

"A=h\nu -E,"

"A=6.63\cdot 10^{-34}\cdot 8\cdot 10^{14}-1.6\cdot 10^{-19}=3.7\cdot 10^{-19}~J."